# Robust $M$-Estimation of Scatter Matrices via Precision Structure Shrinkage

### v1

@@ Abstract
Maronna’s and Tyler’s $M$-estimators are among the most widely used robust estimators for scatter matrices. However, when the dimension of observations is relatively high, their performance can substantially deteriorate in certain situations, particularly in the presence of clustered outliers. To address this issue, we propose an estimator that shrinks the estimated precision matrix toward the identity matrix. We derive a sufficient condition for its existence, discuss its statistical interpretation, and establish upper and lower bounds for its breakdown point. Numerical experiments confirm robustness of the proposed method.
@@ 1 Introduction
Let us consider a $p$-dimensional population distribution with mean vector $\bm{0}$ and covariance matrix $\Sigma$, from which an i.i.d. sample of size $n$ is drawn. The Sample Covariance Matrix (SCM for short) is commonly used as an estimator of a covariance matrix. However, it may perform poorly when the data contain outliers. In fact, the Additive finite-sample Breakdown Point (AdBP for short) of the SCM is $1/(n+1)$, which implies that the inclusion of even one single outlier in the data can severely degrade the performance of the SCM. See Subsection 2.6 for the definition of the AdBP.
@@
To address the presence of outliers, the $M$-estimators of scatter matrices $V$ for elliptical distributions [20], in particular Tyler’s $M$-Estimator (TME for short) that is distribution-free within the class of elliptical distributions [36], have been proposed as classical approaches. These methods simultaneously estimate covariance and precision structures through simple algorithms based on the Mahalanobis distance. Here, throughout this paper, we refer to the inverse of a covariance matrix as the precision matrix, and to scalar multiples of a precision matrix as precision structures. The classical $M$-estimators are not only computationally feasible but also enjoy asymptotic properties such as $\sqrt{n}$-consistency and asymptotic normality [20, 37]. In addition, desirable properties regarding asymptotic relative efficiency and influence functions are also often highlighted. For these reasons, many extensions and variants of these classical $M$-estimators have been studied [1, 5, 8, 9, 13, 16, 23, 24, 32, 34, 35, 40]. However, it is known that the asymptotic breakdown point and AdBP of classical $M$-estimators are at most $1/p$ [9, 16, 20, 37, 39].
@@
Various robust estimators and algorithms have been proposed in order to tackle low breakdown points in high dimensions of classical $M$-estimators. These include the Stahel–Donoho estimator [7, 33, 25], the minimum volume ellipsoid estimator and the minimum covariance determinant estimator [29, 30, 46], the S-estimator [6], the MM-estimator [19], as well as methods based on statistical depth functions [3]. These methods achieve higher breakdown points but are generally more computationally demanding, especially in high-dimensional settings, than the classical $M$-estimators [20, 36]. As explained in [38], a comparison between classical $M$-estimators and these methods from different perspectives is also important. Notably, all the estimators mentioned above, including the classical $M$-estimators, are linearly equivariant. See Remark 2 for the definition of linear equivariance. Advantages of linearly equivariant estimators of $V$ under elliptical distributions lie not only in their explicit equivariance property but also in the applicability of the classical asymptotic theory. In the setting where $p$ is fixed and $n$ tends to infinity, the existence of a probability limit implies the consistency of the estimated covariance structure [21]. On the other hand, for linearly equivariant estimators of $V$, it is known that their replacement finite-sample breakdown points are at most (np+1)/(2n+p1)(n-p+1)/(2n+p-1) when $p<n$ and the data are in general position [6] and so this upper bound is small when $p$ is close to $n$.
@@
To date, various $M$-Estimation methods for $V$, particularly those derived from the TME, have been proposed [6, 8, 13, 32]. Furthermore, penalized $M$-Estimators, in which a penalty term is added to the objective function of the classical $M$-estimation problem, have also been widely studied [23, 34, 40, 44]. Among these approaches, two particularly well-known estimators are the diagonally loaded Normalized Sample Matrix Inversion (LNSMI for short) estimator, also known as the Chen–Wiesel–Hero (CWH for short) estimator, and the Regularized Tyler’s $M$-Estimator (RegTYL for short) [1, 5, 23, 24]. The LNSMI and RegTYL are simple estimators that introduce a ridge-type shrinkage toward the identity matrix $I_{p}$ in the covariance structure estimation part of the TME, and they can be interpreted as incorporating the well-known Ledoit–Wolf (LW- for short) estimator [18] of a covariance matrix [23, 43]. In relation to LNSMI and RegTYL, the Spatial Sign Covariance Matrix (SSCM for short) [41] and the Generalized SSCM (GSSCM for short) have recently been revisited by [27, 40]. Although the SSCM sacrifices linear equivariance, the loss of equivariance enables the estimator to exceed the aforementioned upper bound on the breakdown point [6]. When the population mean is known, the asymptotic breakdown point of SSCM reaches one. Recent research investigates the relationship between estimators including LNSMI and RegTYL, and GSSCM, and examines sufficient conditions under which the AdBP of LNSMI and RegTYL reaches one [40]. These findings suggest that shrinkage toward $I_{p}$ may lead to estimators with strong robustness, yet its effect on robustness has not been thoroughly investigated.
@@
Through numerical experiments, we found that performance of both TME and LNSMI deteriorates in the presence of clustered outliers when $n>p$. The reason is apparently failure of the precision structure. Hence, in this paper, we propose an $M$-estimator of $V$ that introduces ridge-type shrinkage toward $I_{p}$ in a precision structure estimation step. Specifically, we incorporate the Efron–Morris (EM- for short) estimator [11], which is a well-known estimator of the precision matrix along with the Kubokawa–Srivastava (KS- for short) estimator [17], into the $M$-estimation of $V$.
@@
Clustered outliers have been extensively discussed for many years in the context of breakdown-point analysis and numerical studies, and they are regarded as one of the most challenging patterns of outliers in robust estimation of $V$ [2, 22, 31]. Although the MCD and Stahel–Donoho estimators possess high theoretical robustness, several studies have noted practical difficulties under clustered contamination [15, 25, 45]. To address these issues, several approaches have been proposed, which include methods that conduct preprocessing prior to computing a robust estimator and that combine robust estimation with clustering [15, 25, 26, 28, 45]. Our proposed estimator, however, can deal with clustered outliers without requiring such preprocessing, provided appropriately selected shrinkage coefficient.
@@
We examine the theoretical background of the proposed method within the frameworks of penalized $M$-estimation and Bayesian estimation, clarify the conditions for the existence of a solution, and then evaluate the robustness of the proposed method in terms of the AdBP. This study investigates how robustness is affected when a ridge-type shrinkage, which differs from methods such as RegTYL, LNSMI, and the recently proposed Hybrid $M$-Estimator (HME for short) [40], is incorporated into the $M$-Estimation of $V$. Focusing on the framework of penalized $M$-estimation, we numerically demonstrate that the proposed Tyler-type method performs comparably to existing methods such as SCM, TME, LNSMI, SSCM in situations where they operate effectively, while it outperforms them in the presence of clustered outliers. In addition, this paper provides a structured synthesis of the relationships among various classes of estimators within the frameworks of penalized $M$-estimation.
@@
This paper is organized as follows. In Section 2, we precisely formulate the problem and describe the existing methods. In Section 3, we introduce the proposed method, explain its theoretical background, and discuss the conditions for the existence of a solution estimators characterized by estimating equations. In Section 4, we evaluate robustness of the proposed method from the perspective of the AdBP. In Section 5, we examine the performance of the proposed method under various conditions through numerical experiments. The proofs of mathematical statements are given in Section 6. Finally, in Section 7, we summarize the contents of this paper. Details and additional simulation results are provided in the appendix.
@@
$Notations:$ Let $\mathbb{R}_{\geq 0}\coloneqq\{x\in\mathbb{R}\,;\,x\geq 0\}$ and $\mathbb{R}_{>0}\coloneqq\{x\in\mathbb{R}\,;\,x>0\}$. For $I\in\mathbb{N}$, let $I]\coloneqq\{1,\ldots,I\}$. Let $\lceil\cdot\rceil$ denote the ceiling function. Let $(\cdot)^{\top}$ denote the transpose of a matrix or vector. For a nonsingular matrix $A$, let $A^{-1}$ denote its inverse, and let $(A^{-1})^{\top}$ be abbreviated as $A^{-\top}$. For a matrix $A$, let $A^{\dagger}$ denote the Moore–Penrose pseudoinverse of $A$. For a square matrix $A$, let $\mathrm{Tr}[A]$ and $\det(A)$ denote the trace and determinant of $A$, respectively. For a matrix $A$, let $\|A\|_{\textnormal{F}}^{2}\coloneqq\mathrm{Tr}[AA^{\top}]$ denote the squared Frobenius norm of $A$. For a symmetric matrix $A$, let $\lambda_{j}(A)$ denote the $j$-th eigenvalue of $A$. For symmetric matrices $A$ and $B$, let $A\preceq B$ indicate that BAB-A is apositive semidefinite symmetric matrix, which is known as the Löwner partial order. For $\bm{x}\in\mathbb{R}^{p}$, let $\|\bm{x}\|$ denote the Euclidean norm of $\bm{x}$. Let $\mathbb{S}^{p-1}=\{\bm{x}\in\mathbb{R}^{p}\,;\,\|\bm{x}\|=1\}$ denote the unit sphere. Let $\Delta^{p-1}=\{\bm{\eta}\in\mathbb{R}^{p}\,;\,0\leq\eta_{j}\,(j\in[p]),\,\sum_{j=1}^{p}\eta_{j}=1\}$ denote the (p1)(p-1)-dimensional unit simplex. Let $\mathcal{N}_{p}(\bm{0},\Sigma)$ denote the $p$-dimensional normal distribution with mean vector $\bm{0}$ and covariance matrix $\Sigma$. Consider a distribution with mean vector $\bm{0}$ and covariance matrix $\Sigma\in\mathcal{S}_{++}^{p}$. For $\bm{x}\in\mathbb{R}^{p}$, $\bm{x}^{\top}\Sigma^{-1}\bm{x}$ is called the squared Mahalanobis distance of $\bm{x}$, and $\sqrt{\bm{x}^{\top}\Sigma^{-1}\bm{x}}$ is called the Mahalanobis distance of $\bm{x}$ with respect to $\Sigma$. Following the usual convention, for an estimator $\hat{\Sigma}$ of $\Sigma$, $\bm{x}^{\top}\hat{\Sigma}^{-1}\bm{x}$ is also called the squared Mahalanobis distance of $\bm{x}$ with respect to $\hat{\Sigma}$.
@@ 2.1 Setup and notation for the probability space
Hereafter, we fix a single measurable space $(\Xi,\mathcal{F})$. Consider a probability space $(\Xi,\mathcal{F},\mathcal{P})$ equipped with a probability measure $\mathcal{P}$ and a general measurable space $(S,\mathcal{A})$, and let $\bm{R}:(\Xi,\mathcal{F})\to(S,\mathcal{A})$ be a random vector or random variable. Let us denote by $P_{\bm{R}}(\cdot)$ the measure on $(S,\mathcal{A})$ induced by $\bm{R}$, that is, $P_{\bm{R}}(\cdot)\coloneqq\mathcal{P}(\bm{R}^{-1}(\cdot))$. Moreover, for $\chi=\mathbb{R}^{p}$, we consider the observation space $(\chi,\mathcal{B}(\chi))$. For a random vector $\bm{X}:(\Xi,\mathcal{F})\to(\chi,\mathcal{B}(\chi))$, we write $P_{\bm{X}}(\cdot)$ simply as $P(\cdot)$ when no confusion arises from the context, and define probabilities by $P(\text{condition})\coloneqq P_{\bm{X}}({\bm{X}\in\chi;\text{condition}})$.
@@ 2.2 Elliptical distributions
we say that $\bm{X}$ follows a (centered) elliptical distribution, and denote it by $\bm{X}\sim\textnormal{ES}_{p}(\bm{0},V,g(\cdot))$. The function, $g:\mathbb{R}_{\geq 0}\to\mathbb{R}_{>0}$ is a nonincreasing function, referred to as the density generator. Moreover, $C_{p,g}>0$ is a normalizing constant depending on $p$ and $g(\cdot)$, and $V\in\mathcal{S}_{++}^{p}$ is referred to as the scatter matrix or shape matrix. An elliptical distribution is characterized by the property that the contours of its PDF are ellipsoidal. For example, the multivariate normal, the multivariate $t$, and the multivariate Cauchy distributions are included in the class of the elliptical distributions. For an elliptical distribution with a finite covariance matrix $\Sigma\in\mathcal{S}_{++}^{p}$, there exists a constant $c\in\mathbb{R}_{>0}$ such that $\Sigma=cV$. Therefore, the scatter matrix $V$ is also referred to as the covariance structure.
2.3 Formulation of problem
@@
We consider the case where the mean is known and an i.i.d. sample $\bm{X}_{1},\ldots,\bm{X}_{n}$ of size $n$ is drawn from the population distribution $\textnormal{ES}_{p}(\bm{0},V;g(\cdot))$. We assume that $n>p$, and treat the density generator $g(\cdot)$ as an unknown but fixed function. Let $\bm{x}_{1},\ldots,\bm{x}_{n}$ denote realizations of $\bm{X}_{1},\ldots,\bm{X}_{n}$. We then consider the problem of estimating the scatter matrix $V\in\Theta$ based on the observed data $\bm{x}_{1},\ldots,\bm{x}_{n}$}.

### v2

@@ Abstract
Maronna’s and Tyler’s $M$-estimators are among the most widely used robust estimators for scatter matrices. However, when the dimension of observations is relatively high, their performance can substantially deteriorate in certain situations, particularly in the presence of clustered outliers. To address this issue, we propose an estimator that shrinks the estimated precision matrix toward the identity matrix. We derive a sufficient condition for its existence, discuss its statistical interpretation, and establish upper and lower bounds for its additive finite sample breakdown point. Numerical experiments confirm the robustness of the proposed method.
@@ 1 Introduction
Let us consider a $p$-dimensional population distribution with mean vector $\bm{0}$ and covariance matrix $\Sigma$, from which an i.i.d. sample of size $n$ is drawn. The Sample Covariance Matrix (SCM for short) is commonly used as an estimator of a covariance matrix. However, it may perform poorly when the data contain outliers. In fact, the Additive finite sample Breakdown Point (AdBP for short) of the SCM is $1/(n+1)$, which implies that the inclusion of even one single outlier in the data can severely degrade the performance of the SCM. See Subsection 2.6 for the definition of the AdBP.
@@
To address the presence of outliers, the $M$-estimators of scatter matrices $V$ for elliptical distributions [19], in particular Tyler’s $M$-Estimator (TME for short) that is distribution-free within the class of elliptical distributions [35], have been proposed as classical approaches. These methods simultaneously estimate covariance and precision structures through simple algorithms based on the Mahalanobis distance. Here, throughout this paper, we refer to the inverse of a covariance matrix as the precision matrix and to scalar multiples of a precision matrix as precision structures. The classical $M$-estimators are not only computationally feasible but also enjoy asymptotic properties such as $\sqrt{n}$-consistency and asymptotic normality [19, 36]. In addition, desirable properties regarding asymptotic relative efficiency and influence functions are also often highlighted. For these reasons, many extensions and variants of these classical $M$-estimators have been studied [1, 5, 7, 8, 12, 22, 30, 33, 39]. However, it is known that the asymptotic breakdown point and AdBP of classical $M$-estimators are at most $1/p$ [8, 19, 38].
@@
Various robust estimators and algorithms have been proposed in order to tackle low breakdown points of classical $M$-estimators in high dimensional settings. These include the Stahel–Donoho (SD- for short) estimator [31], the minimum volume ellipsoid estimator and the Minimum Covariance Determinant (MCD- for short) estimator [28], the S-estimator [6], the MM-estimator [34], as well as methods based on statistical depth functions [3]. These methods achieve higher breakdown points but are generally more computationally demanding, especially in high dimensional settings, than the classical $M$-estimators [19, 35]. A comparison between classical $M$-estimators and these methods from different perspectives remains important [37], even though algorithmic and implementation level efforts to improve their computational efficiency have been reported [24, 25, 29, 45]. Notably, all the estimators mentioned above, including the classical $M$-estimators, are linearly equivariant. See Remark 2 for the definition of linear equivariance. Advantages of linearly equivariant estimators of $V$ under elliptical distributions lie not only in their explicit equivariance property but also in the applicability of the classical asymptotic theory. In the setting where $p$ is fixed and $n$ tends to infinity, the existence of a probability limit implies the consistency of the estimated covariance structure [20]. On the other hand, for linearly equivariant estimators of $V$, it is known that their replacement finite sample breakdown points are at most (np+1)/(2n+p1)(n-p+1)/(2n+p-1) when $p<n$ and the data are in general position [6] and so this upper bound is small when $p$ is close to $n$.
@@
To date, various $M$-estimation methods for $V$, particularly those derived from the TME, have been proposed [7, 12, 30]. Furthermore, penalized $M$-estimators, in which a penalty term is added to the objective function of the classical $M$-estimation problem, have also been widely studied [22, 23, 32, 39, 43, 46]. Among these approaches, two particularly well-known estimators are the diagonally loaded Normalized Sample Matrix Inversion (LNSMI for short) estimator, also known as the Chen–Wiesel–Hero (CWH for short) estimator, and the Regularized Tyler’s $M$-estimator (RegTYL for short) [1, 5, 22]. The LNSMI and RegTYL are simple estimators that introduce a ridge type shrinkage toward the identity matrix $I_{p}$ in the covariance structure estimation part of the TME, and they can be interpreted as incorporating the well-known Ledoit–Wolf (LW- for short) estimator [18] of a covariance matrix [22, 42]. In relation to LNSMI and RegTYL, the Spatial Sign Covariance Matrix (SSCM for short) [40] and the Generalized SSCM (GSSCM for short) have recently been revisited by [26, 39]. Although the SSCM sacrifices linear equivariance, the loss of equivariance enables the estimator to exceed the aforementioned upper bound on the breakdown point [6]. When the population mean is known, the asymptotic breakdown point of SSCM reaches one. Recent research investigates the relationship between estimators including LNSMI and RegTYL, and GSSCM, and examines sufficient conditions under which the AdBP of LNSMI and RegTYL reaches one [39]. These findings suggest that shrinkage toward $I_{p}$ may lead to estimators with strong robustness, yet its effect on robustness has not been thoroughly investigated.
@@
Through numerical experiments, we found that the performance of both TME and LNSMI deteriorates in the presence of clustered outliers when $n>p$. The reason appears to be a failure of the precision structure. Hence, in this paper, we propose an $M$-estimator of $V$ that introduces ridge type shrinkage toward $I_{p}$ in a precision structure estimation step. Specifically, we incorporate the Efron–Morris (EM- for short) estimator [10], which is a well-known estimator of the precision matrix along with the Kubokawa–Srivastava (KS- for short) estimator [17], into the $M$-estimation of $V$.
@@
Clustered outliers have been extensively discussed for many years in the context of breakdown point analysis and numerical studies, and they are regarded as one of the most challenging patterns of outliers for robust estimation of $V$ [2, 21, 38]. Although the MCD and SD-estimators possess high theoretical robustness, clustered contamination has been discussed in several applied and computational studies [14, 15, 24, 25, 44, 27]. To address these issues, several approaches have been proposed [24, 25], which include methods that conduct preprocessing prior to computing a robust estimator [15]. Our proposed estimator, however, can deal with clustered outliers without requiring such preprocessing, provided an appropriately selected shrinkage coefficient.
@@
We examine the theoretical background of the proposed estimator within the frameworks of penalized $M$-estimation and Bayesian estimation, clarify the conditions for the existence of a solution, and then evaluate the robustness of the proposed estimator in terms of the AdBP. This study investigates how robustness is affected when a ridge type shrinkage, which differs from those used in methods such as RegTYL, LNSMI, and the recently proposed Hybrid $M$-Estimator (HME for short) [39], is incorporated into the $M$-estimation of $V$. Focusing on the framework of penalized $M$-estimation, we numerically demonstrate that the proposed method performs comparably to existing methods such as SCM, TME, LNSMI, and SSCM in situations where they operate effectively, while it outperforms them in the presence of clustered outliers. In addition, this paper provides a structured synthesis of the relationships among various classes of estimators within the frameworks of penalized $M$-estimation.
@@
This paper is organized as follows. In Section 2, we precisely formulate the problem and describe the existing estimators. In Section 3, we introduce the proposed estimator, explain its theoretical background, and discuss the conditions for the existence of estimators characterized by estimating equations. In Section 4, we evaluate the robustness of the proposed estimator from the perspective of the AdBP. In Section 5, we examine the performance of the proposed estimator under various conditions through numerical experiments. The proofs of mathematical statements are given in Section 6. Finally, in Section 7, we summarize the contents of this paper. Details and additional simulation results are provided in the Appendix.
@@
Notations: Let $\mathbb{R}_{\geq 0}\coloneqq\{x\in\mathbb{R}:x\geq 0\}$ and $\mathbb{R}_{>0}\coloneqq\{x\in\mathbb{R}:x>0\}$. For $I\in\mathbb{N}$, let $I]\coloneqq\{1,\ldots,I\}$. Let $\lceil\cdot\rceil$ denote the ceiling function. Let $(\cdot)^{\top}$ denote the transpose of a matrix or vector. For a nonsingular matrix $A$, let $A^{-1}$ denote its inverse, and let $(A^{-1})^{\top}$ be abbreviated as $A^{-\top}$. For a matrix $A$, let $A^{\dagger}$ denote the Moore–Penrose pseudoinverse of $A$. For a square matrix $A$, let $\mathrm{Tr}[A]$ and $\det(A)$ denote the trace and determinant of $A$, respectively. For a matrix $A$, let $\|A\|_{\mathrm{F}}^{2}\coloneqq\mathrm{Tr}[AA^{\top}]$ denote the squared Frobenius norm of $A$. For a symmetric matrix $A$, let $\lambda_{j}(A)$ denote the $j$-th eigenvalue of $A$. For symmetric matrices $A$ and $B$, let $A\preceq B$ indicate that BAB-A is a positive semidefinite symmetric matrix, which is known as the Löwner partial order. For $\bm{x}\in\mathbb{R}^{p}$, let $\|\bm{x}\|$ denote the Euclidean norm of $\bm{x}$. Let $\mathbb{S}^{p-1}=\{\bm{x}\in\mathbb{R}^{p}:\|\bm{x}\|=1\}$ denote the unit sphere. Let $\Delta^{p-1}=\{\bm{\eta}\in\mathbb{R}^{p}:0\leq\eta_{j}\,(j\in[p]),\,\sum_{j=1}^{p}\eta_{j}=1\}$ denote the (p1)(p-1)-dimensional unit simplex. Let $\mathcal{N}_{p}(\bm{0},\Sigma)$ denote the $p$-dimensional normal distribution with mean vector $\bm{0}$ and covariance matrix $\Sigma$. Consider a distribution with mean vector $\bm{0}$ and covariance matrix $\Sigma\in\mathcal{S}_{++}^{p}$. For $\bm{x}\in\mathbb{R}^{p}$, $\bm{x}^{\top}\Sigma^{-1}\bm{x}$ is called the squared Mahalanobis distance of $\bm{x}$, and $\sqrt{\bm{x}^{\top}\Sigma^{-1}\bm{x}}$ is called the Mahalanobis distance of $\bm{x}$ with respect to $\Sigma$. Following the usual convention, for an estimator $\hat{\Sigma}$ of $\Sigma$, $\bm{x}^{\top}\hat{\Sigma}^{-1}\bm{x}$ is also called the squared Mahalanobis distance of $\bm{x}$ with respect to $\hat{\Sigma}$.
@@ 2.1 Setup and notation for the probability space
Hereafter, we fix a single measurable space $(\Xi,\mathcal{F})$. Consider a probability space $(\Xi,\mathcal{F},\mathcal{P})$ equipped with a probability measure $\mathcal{P}$ and a general measurable space $(S,\mathcal{A})$, and let $\bm{R}:(\Xi,\mathcal{F})\to(S,\mathcal{A})$ be a random vector or random variable. Let us denote by $P_{\bm{R}}(\cdot)$ the measure on $(S,\mathcal{A})$ induced by $\bm{R}$, that is, $P_{\bm{R}}(\cdot)\coloneqq\mathcal{P}(\bm{R}^{-1}(\cdot))$. Moreover, for $\chi=\mathbb{R}^{p}$, we consider the observation space $(\chi,\mathcal{B}(\chi))$. For a random vector $\bm{X}:(\Xi,\mathcal{F})\to(\chi,\mathcal{B}(\chi))$, we write $P_{\bm{X}}(\cdot)$ simply as $P(\cdot)$ when no confusion arises from the context, and define probabilities by $P(\text{condition})\coloneqq P_{\bm{X}}({\bm{X}\in\chi:\text{condition}})$.
@@ 2.2 Elliptical distributions
we say that $\bm{X}$ follows a (centered) elliptical distribution, and denote it by $\bm{X}\sim\textnormal{ES}_{p}(\bm{0},V,g(\cdot))$. The function, $g:\mathbb{R}_{\geq 0}\to\mathbb{R}_{>0}$ is a nonincreasing function, referred to as the density generator. Moreover, $C_{p,g}>0$ is a normalizing constant depending on $p$ and $g(\cdot)$, and $V\in\mathcal{S}_{++}^{p}$ is referred to as the scatter matrix or shape matrix. An elliptical distribution is characterized by the property that the contours of its PDF are ellipsoidal. For example, the multivariate normal, the multivariate $t$, and the multivariate Cauchy distributions are included in the class of elliptical distributions. For an elliptical distribution with a finite covariance matrix $\Sigma\in\mathcal{S}_{++}^{p}$, there exists a constant $c\in\mathbb{R}_{>0}$ such that $\Sigma=cV$. Therefore, the scatter matrix $V$ is also referred to as the covariance structure.
2.3 Formulation of the problem
@@
We consider the case where the mean is known and an i.i.d. sample $\bm{X}_{1},\ldots,\bm{X}_{n}$ of size $n$ is drawn from the population distribution $\textnormal{ES}_{p}(\bm{0},V,g(\cdot))$. We assume that $n>p$, and treat the density generator $g(\cdot)$ as an unknown but fixed function. Let $\bm{x}_{1},\ldots,\bm{x}_{n}$ denote realizations of $\bm{X}_{1},\ldots,\bm{X}_{n}$. We then consider the problem of estimating the scatter matrix $V\in\Theta$ based on the observed data $\bm{x}_{1},\ldots,\bm{x}_{n}$}.